# Generalized Lie Symmetries in Physics-Informed Neural Operators

### v1

@@ 1 Introduction
Deep neural networks are increasingly used for scientific computing, particularly in simulating complex physical systems governed by partial differential equations (PDEs) [11], where traditional methods struggle [24].
Physics-informed neural networks (PINNs) [24] integrate physical laws and constraints directly into the learning process, demonstrating success in various applications [24, 29, 35, 14, 27, 33]. However, classical PINNs face challenges under varying parameters and boundary conditions and often encounter training challenges due to unbalanced gradients and non-convex loss landscapes [10, 15, 22]. Neural operators offer a compelling alternative, which unlike standard feedforward neural networks, learn mappings between infinite-dimensional function spaces, directly connecting parameters and initial/boundary conditions to PDE solutions [21, 32]. This eliminates the need for independent simulations and, when combined with physics-informed loss functions, often improves generalization and physical validity. We focus on this class of approaches.
In a separate direction, the field of geometric deep learning has demonstrated the practical and theoretical benefits of incorporating symmetry information into neural networks, leading to improved performance through beneficial inductive biases [6, 5]. For example, the translational equivariance of convolutional neural networks (CNNs) [9, 18] has been suggested as an important reason for their successes in tasks such as image classification [16, 12].
Similarly, recent work in neural operators has shown that incorporating PDE symmetries improves generalization and training efficiency. However, most research has focused on Lie point symmetries due to their well-established theoretical foundation and systematic derivation methods [23, 13, 3]. In contrast, the potential of generalized symmetries remains largely unexplored due to their unsuitability for data augmentation or equivariant architectures.
@@ 1.1 PDE Symmetries in Deep Learning
is the simplest way to add symmetry information into neural PDE solvers [4, 20] by augmenting the training data with transformed versions obtained by applying symmetry operations. While straightforward to implement, this approach can increase training costs.
@@
is an alternative way to inject symmetry information by appropriately regularizing the loss using symmetry information [1, 19, 34]. This often leads to improved generalization and sample efficiency, although it may not guarantee equivariant models.
@@
aims to directly embed symmetries into the model architecture. Given the complexity of point symmetry groups associated with PDEs, this requires either general techniques [26] or exploiting simpler subgroups with existing architectures [31].
Beyond these, alternative approaches exist, based on problem reformulations [2, 17] or those focused on other types of structures, e.g. from Hamiltonian dynamics [7, 8, 30].
@@ 2 PDE Symmetries
Here, we introduce the notion of PDE symmetries. The formal presentation here closely follows [23], and the reader is strongly encouraged to read it for more in-depth discussions of the concepts.
@@
However, it is important to note that these are not symmetries in the traditional sense, as they no longer act on $X\times U$. For a more in-depth discussion, see [23].
@@ 3 Methodology
Table 1: Comparison of $L_{2}$ loss and equation error for Darcy Flow and Burgers’ equation with varying sample sizes, using our symmetry-based method, PINO (no symmetry), and the method from [1, 34].
Equation No. Samples Metric Symmetry (Ours) No Symmetry (PINO) Symmetry [1, 34]
@@
To understand the proposed methodology, we first recap in simpler notation, the method of [1, 34]. Clasically, PINNs aim to solve the PDE by minimizing the residual
@@
In [1, 34], it is proposed that for a Lie algebra $v_{i}$, it is beneficial to minimize (for some $\gamma>0$):
@@ 4.1 Darcy Flow
on the domain $x\in[0,1]^{2}$ with $u(x)=0$ on the boundary, where $k(x)$ is the permeability field and $f(x)$ is the source term. We tested our approach on a $61\times 61$ resolution dataset downsampled from a $421\times 421$ resolution dataset and set $f({x})=1$ as in [21]. We trained a 2D Fourier neural operator model for 300 epochs using the proposed generalized symmetry loss in Equation 3. As is common with PINOs, in this example we also include a data loss term enforcing initial and boundary conditions, and a trajectory matching loss.The trained model is tested on 500 samples. Table 1 indicates that the proposed symmetry regularization improves prediction in terms of both $L_{2}$ and equation losses, being more sample efficient in minimizing the PDE residuals. As shown in Equation 10, the point symmetry method of [1, 34] in Equation 2 generates no extra terms, being equivalent to a standard PINO.
@@ 4.2 Burgers’ equation
on the domain $x\in[0,1]$ and $t\in[0,1]$ with initial condition $u(x,0)=u_{0}(x)$, for which we generated the 1D Burgers’ dataset following [25], using a $128\times 100$ grid and setting the viscosity coefficient to $\nu=0.01$. We compared the proposed method in Equation 3 with [1, 34] defined in Equation 2. Table 1 presents the results, demonstrating that the proposed method consistently achieves either superior or comparable performance to the baselines with respect to both equation error and $L_{2}$ error. Following experimental details of [1], the residual for this equation does not appear directly in the training objective. Figure 2 provides a representative prediction showcasing the improved error achieved by our method.
@@ 5 Conclusion
We further posit that the observed efficiency partly arises from the inducement of a Sobolev-type norm [28]. Crucially, our approach offers a systematic way for selecting the appropriate norm based on the governing PDE. This connection between generalized symmetries, Sobolev norms, and improved training warrants further investigation.

### v2

@@ 1 Introduction
Deep neural networks are increasingly used for scientific computing, particularly in simulating complex physical systems governed by partial differential equations (PDEs) han2018solving , where traditional methods struggle raissi2019physics .
Physics-informed neural networks (PINNs) (raissi2019physics, ) integrate physical laws and constraints directly into the learning process, demonstrating success in various applications (raissi2019physics, ; sun2020surrogate, ; zhu2019physics, ; karumuri2020simulator, ; sirignano2018dgm, ; zhang2024biophysics, ). However, classical PINNs face challenges under varying parameters and boundary conditions and often encounter training challenges due to unbalanced gradients and non-convex loss landscapes (grossmann2023can, ; krishnapriyan2021characterizing, ; mullerposition, ). Neural operators offer a compelling alternative, which unlike standard feedforward neural networks, learn mappings between infinite-dimensional function spaces, directly connecting parameters and initial/boundary conditions to PDE solutions li2021physics ; wang2021learning . This eliminates the need for independent simulations and, when combined with physics-informed loss functions, often improves generalization and physical validity. We focus on this class of approaches.
In a separate direction, the field of geometric deep learning has demonstrated the practical and theoretical benefits of incorporating symmetry information into neural networks, leading to improved performance through beneficial inductive biases (bronstein2021geometric, ; brehmer2024doesequivariancematterscale, ). For example, the translational equivariance of convolutional neural networks (CNNs) (fukushima1982neocognitron, ; lecun1989backpropagation, ) has been suggested as an important reason for their successes in tasks such as image classification (krizhevsky2012imagenet, ; he2016deep, ).
Similarly, recent work in neural operators has shown that incorporating PDE symmetries improves generalization and training efficiency. However, most research has focused on Lie point symmetries due to their well-established theoretical foundation and systematic derivation methods (olver1993applications, ; ibragimov1995crc, ; Baumann1998MathLieAP, ). In contrast, the potential of generalized symmetries remains largely unexplored due to their unsuitability for data augmentation or equivariant architectures.
@@ 1.1 PDE Symmetries in Deep Learning
is the simplest way to add symmetry information into neural PDE solvers (brandstetter2022lie, ; 10016380, ) by augmenting the training data with transformed versions obtained by applying symmetry operations. While straightforward to implement, this approach can increase training costs.
@@
is an alternative way to inject symmetry information by appropriately regularizing the loss using symmetry information (akhoundsadegh2023lie, ; li2023utilizing, ; zhang2023enforcing, ). This often leads to improved generalization and sample efficiency, although it may not guarantee equivariant models.
@@
aims to directly embed symmetries into the model architecture. Given the complexity of point symmetry groups associated with PDEs, this requires either general techniques (shumaylov2024lie, ) or exploiting simpler subgroups with existing architectures (wang2021incorporating, ).
Beyond these, alternative approaches exist, based on problem reformulations (arora2024invariant, ; lagrave2022equivariant, ) or those focused on other types of structures, e.g. from Hamiltonian dynamics (canizares2024hamiltonian, ; canizares2024symplectic, ; tanaka2024neuraloperatorsmeetenergybased, ).
@@ 2 PDE Symmetries
Here, we introduce the notion of PDE symmetries. The formal presentation here closely follows (olver1993applications, ), and the reader is strongly encouraged to read it for more in-depth discussions of the concepts.
@@
However, it is important to note that these are not symmetries in the traditional sense, as they no longer act on $X\times U$. For a more in-depth discussion, see olver1993applications .
@@ 3 Methodology
Table 1: Comparison of $L_{2}$ loss and equation error for Darcy Flow and Burgers’ equation with varying sample sizes, using our symmetry-based method, PINO (no symmetry), and the method from akhoundsadegh2023lie ; zhang2023enforcing .
Equation No. Samples Metric Symmetry (Ours) No Symmetry (PINO) Symmetry akhoundsadegh2023lie ; zhang2023enforcing
@@
To understand the proposed methodology, we first recap in simpler notation, the method of (akhoundsadegh2023lie, ; zhang2023enforcing, ). Clasically, PINNs aim to solve the PDE by minimizing the residual
@@
In (akhoundsadegh2023lie, ; zhang2023enforcing, ), it is proposed that for a Lie algebra $v_{i}$, it is beneficial to minimize (for some $\gamma>0$):
@@ 4.1 Darcy Flow
on the domain $x\in[0,1]^{2}$ with $u(x)=0$ on the boundary, where $k(x)$ is the permeability field and $f(x)$ is the source term. We tested our approach on a $61\times 61$ resolution dataset downsampled from a $421\times 421$ resolution dataset and set $f({x})=1$ as in li2021physics . We trained a 2D Fourier neural operator model for 300 epochs using the proposed generalized symmetry loss in Equation 3. As is common with PINOs, in this example we also include a data loss term enforcing initial and boundary conditions, and a trajectory matching loss. The trained model is tested on 500 samples. Table 1 indicates that the proposed symmetry regularization improves prediction in terms of both $L_{2}$ and equation losses, being more sample efficient in minimizing the PDE residuals. As shown in Equation 10, the point symmetry method of akhoundsadegh2023lie ; zhang2023enforcing in Equation 2 generates no extra terms, being equivalent to a standard PINO.
@@ 4.2 Burgers’ equation
on the domain $x\in[0,1]$ and $t\in[0,1]$ with initial condition $u(x,0)=u_{0}(x)$, for which we generated the 1D Burgers’ dataset following rosofsky2023applications , using a $128\times 100$ grid and setting the viscosity coefficient to $\nu=0.01$. We compared the proposed method in Equation 3 with akhoundsadegh2023lie ; zhang2023enforcing defined in Equation 2. Table 1 presents the results, demonstrating that the proposed method consistently achieves either superior or comparable performance to the baselines with respect to both equation error and $L_{2}$ error. Following experimental details of akhoundsadegh2023lie , the residual for this equation does not appear directly in the training objective. Figure 2 provides a representative prediction showcasing the improved error achieved by our method.
@@
4.3 Ablation studies
We evaluated the robustness, generalization ability, and stability of our approach and PINO by adding various levels of noise into the training dataset using the Darcy flow equation with 500 samples, as shown in Table 2. The generalization gap, calculated as the difference between the $L_{2}$ loss on the test and training sets, increases slightly with higher noise level. This suggests that our approach generalizes reasonably well and resists overfitting to noisy data.
For robustness, both test $L_{2}$ loss and the equation error show a steady increase as the noise level rises, indicating that the model remains relatively robust up to $5\%$ noise level. In comparison, our approach consistently achieves lower $L_{2}$ loss and the equation error than PINO as the noise level increase.
Table 2: Ablation study of model robustness, generalization, and stability using Darcy flow
Metric Clean data ($0\%$) Noise ($1\%$) Noise ($5\%$) Noise ($10\%$)
Ours PINO Ours PINO Ours PINO Ours PINO
$L_{2}$ Loss (Train) $\bm{0.007\pm 6e-6}$ $60.007\pm 6e-$ 6 $60.007\pm 6e-$ 6 $60.007\pm 6e-$ 6 $60.008\pm 6e-$ 6 $60.008\pm 6e-$ 6 $60.010\pm 4e-$ 6 $60.010\pm 5e-$ 6
$L_{2}$ Loss (Test) $\bm{0.016\pm 4e-4}$ $40.019\pm 5e-$ 4 $40.018\pm 4e-$ 4 $40.019\pm 5e-$ 4 $40.019\pm 4e-$ 4 $40.020\pm 4e-$ 4 $40.022\pm 4e-$ 4 $40.022\pm 5e-$ 4
Eqn. error (Train) $\bm{0.007\pm 6e-4}$ $40.007\pm 6e-$ 4 $40.092\pm 5e-$ 4 $40.091\pm 5e-$ 4 $30.264\pm 1e-$ 3 $30.268\pm 1e-$ 3 $30.461\pm 2e-$ 3 $30.488\pm 2e-$ 3
Eqn. error (Test) $30.569\pm 7e-$ 3 $30.701\pm 9e-$ 3 $30.624\pm 8e-$ 3 $20.723\pm 9e-$ 2 $30.722\pm 8e-$ 3 $30.723\pm 9e-$ 3 $30.827\pm 4e-$ 3 $30.918\pm 9e-$ 3
Generalization gap $\bm{0.009\pm 4e-4}$ $40.012\pm 5e-$ 4 $40.011\pm 4e-$ 4 $40.012\pm 5e-$ 4 $40.011\pm 4e-$ 4 $40.012\pm 4e-$ 4 $40.012\pm 4e-$ 4 $40.012\pm 5e-$ 4
Stability gap $30.562\pm 7e-$ 3 $30.694\pm 9e-$ 3 $30.532\pm 8e-$ 3 $30.632\pm 9e-$ 3 $30.458\pm 8e-$ 3 $30.455\pm 9e-$ 3 $\bm{0.366\pm 4e-3}$ $30.430\pm 9e-$ 3
However, noise level does impact the the model’s ability to capture the underlying physics well as reflected by the stability gap, defined as the difference between the equation errors on the test and training sets. Notably, while both the training and test equation errors increase with noise, the test equation error grows more slowly. Compared to PINO, our approach achieves a lower stability gap by having smaller equation errors, indicating that incorporating symmetry enhances its ability to learn the underlying physical laws.
@@ 5 Conclusion
We further posit that the observed efficiency partly arises from the inducement of a Sobolev-type norm son2023sobolev . Crucially, our approach offers a systematic way for selecting the appropriate norm based on the governing PDE. This connection between generalized symmetries, Sobolev norms, and improved training warrants further investigation.